# Viewpoint Consistency in 3D Generation via Training-free Attention and CLIP Guidance

### v3

@@ 1 Introduction
Figure 2: The long-tailed distribution of viewpoints in the diffusion model’s training data biases the generated viewpoints. The left figure shows a random sample of 6,762 images from Laion2B [31], categorized into three object viewpoints using CLIP ViT-H/14 [13]. The diffusion model maps this long-tailed distribution to a Gaussian distribution, which then produces a similarly biased generation distribution via reverse SDE—fundamentally causing incorrect guidance in 3D representations in the Janus Problem.
Previous studies have attributed the Janus Problem to a generative bias inherent in diffusion models [1, 12, 10, 21]. Given an input prompt, diffusion models tend to generate front-facing images of objects, contributing to geometric inconsistencies in the resulting 3D models. Some methods attempt to mitigate this issue by fine-tuning diffusion models on specialized 3D datasets to enhance multi-view generation capabilities [21, 34]. However, directly fine-tuning on a 3D dataset can lead to catastrophic forgetting and overfitting [11, 30, 42], in addition to being time-consuming and computationally intensive. Alternatively, other approaches address the issue by adjusting sampling strategies to achieve more balanced samples [12]. Nevertheless, without precise knowledge of the long-tailed distribution of the data, such strategies rely on diffusion models to estimate the distribution during sampling, which is often inaccurate [43].
@@ 2.1 Text-to-3D Generation
This task of generating 3D content from text can be broadly divided into two categories. The first category includes end-to-end models such as Point-E [25] and Shap-E [15], which generate 3D outputs directly from text. These models offer fast generation but suffer from poor generalization and subpar quality due to the limited availability of 3D datasets [20]. The second category follows a 2D-to-3D approach inspired by DreamFields [14], where CLIP [27] supervises NeRF [24, 41]. However, the limited supervisory effectiveness of CLIP has prevented this approach from gaining widespread adoption. DreamFusion [26] introduced Score Distillation Sampling (SDS), leveraging pre-trained diffusion models to guide NeRF optimization, which inspired numerous follow-up studies.
Subsequent improvements to DreamFusion have focused on three areas: generation quality, speed, and the Janus problem [40, 19, 18, 4, 12, 21, 5]. Efforts to improve quality mainly revolve around optimizing SDS, as seen in works like ProlificDreamer [40] and LucidDreamer [18], which analyze how SDS leads to mode collapse and excessive smoothing. Speed improvements typically involve optimizing the 3D representation and the generation framework. For example, Magic3D [19] introduces a two-stage optimization strategy and incorporates DMTet [33] for fast surface optimization, while DreamGaussian [39] employs 3D Gaussian Splatting [16] to accelerate generation. The Janus problem, which concerns generating consistent front and back views, is the primary focus of our research. Previous solutions rely on fine-tuning diffusion models to obtain multi-view capabilities [21, 34]. In contrast, our work takes a different approach by analyzing the inherent generative bias of diffusion models. We achieve superior results without compromising the model’s generalization ability.
@@ 3.1 Rethink Janus Problem
Although numerous studies have identified generation bias in diffusion models as a cause of the Janus Problem [21, 1, 10, 12], a comprehensive proof is still lacking, leaving the issue unresolved. This chapter aims to demonstrate that the long-tail distribution of viewpoints in the training data of diffusion models is the fundamental factor inducing generation bias, and that the resulting imbalance in viewpoint category guidance for 3D models constitutes the essence of the Janus Problem.
Figure 4: Overview of our work. We adjust cross-attention maps in the diffusion model to enhance focus on viewpoint keywords, guiding the model to generate diverse viewpoints. CLIP guidance prunes incorrect viewpoint during denoising based on a loss threshold, ensuring alignment with specified viewpoints. Decompose the original prompt into object and descriptive words, and optimize the object prompt in the early stages to ensure geometric integrity.
@@
which describes the evolution of the Gaussian distribution over time $s$ from 0 to $T$, corresponding to $t$ decreasing from $T$ to 0. This is essentially the reverse of Eq. 5. See supplementary for details.
@@
In this context, we interpret SDS as a mechanism that use images generated from diffusion models [18] to guide images rendered by 3D model. The long-tailed distribution of viewpoints in the training data—where the majority of images are front-facing—introduces bias into the generated images [32], as shown in Figure 2. When the generated images exhibit a similarly long-tailed distribution, the guidance becomes imbalanced, causing the 3D object to overfit to images from the highest-density region.
@@ 3.2 Cross-attention Control
During each noise prediction step in the U-Net of the diffusion process, we identify viewpoint-related keywords in the text prompt—terms that specify spatial viewpoint such as “front”, “side”, or “back”. These keywords are associated with spatial information that influences the viewpoint of the generated image. We dynamically adjust the attention weights of these viewpoint keywords based on the desired rendering perspective. This adjustment is implemented using the following formula:
@@
Figure 6: Comparison with baselines methods in text-to-3D generation. Adding our method to the baseline significantly improves geometric consistency without increasing optimization time or compromising generation quality.
@@ 3.3 CLIP Pruning
where $e_{\text{image}}$ and $e_{\text{text}}$ represent the embeddings of $\bar{x}_{0}$ and $P_{v}$, respectively. The dot product measures semantic alignment, and $\sigma$ is normalized to a range between 1-1 and $1$. If $\sigma$ falls below a specified threshold $\tau$, it indicates that the current generation provides incorrect guidance for the 3D model under the specified viewpoint In such cases, the update step is skipped to prevent the inclusion of misleading guidance that could cause the Janus problem.
@@
CLIP Pruning mitigates generation bias by eliminating guidance with insufficient semantic view alignment, thereby balancing the distribution of pseudo-GT viewpoints during training. This pruning strategy rebalances the generation bias band to some extent, leading to a uniform sampling distribution of viewpoints in 3D generation.
@@ 3.4 Coarse-to-Fine Optimization
We first divide the input prompt into an object prompt (main subject) and a description prompt (additional characteristics). In the initial stage, we focus on the object prompt by adjusting cross-attention maps and applying CLIP pruning to amplify attention to spatial viewpint terms. This ensures the generated 3D model accurately captures the core geometry and specified viewpoint.
@@ 4.1 Text-to-3D Generation
Quantitative Comparison. Previous research in 3D generation lacks standardized metrics. Existing works typically use 2D image evaluation metrics to assess the quality or prompt relevance of generated content. However, these metrics are insufficient for evaluating the Janus problem. Therefore, we introduce a novel metric called View-Dependent CLIP-Score.
We generate 3D objects using 20 randomly selected text prompts from the DreamFusion prompt dataset. For each 3D object, we render four viewpoints with uniformly distributed azimuth angles. The View-CS score is calculated by appending a ”* view” description to each input text prompt and computing the similarity between this modified prompt and the corresponding rendered image. The average similarity score across all views constitutes the final metric.
Additionally, to assess the consistency of 3D geometry, we count the instances of inconsistent 3D content produced by each approach, defining this count as the Janus Problem Rate (JR). As shown in Table 1, our method significantly reduces the multi-head problem and improves the View-CS.
@@
LucidDreamer [18] 80.00 29.40
LucidDreamer+ACG (Ours) 30.00 30.26
Magic3D [19] 75.00 28.20
Magic3D+ACG (Ours) 35.00 29.98
DreamFusion [26] 80.00 29.31
DreamFusion+ACG (Ours) 35.00 30.13
Table 1: Comparison of geometric consistency between baseline methods and ACG-enhanced versions. Lower JR indicates fewer Janus outcomes, while higher View-CS reflects better alignment with view-specific prompts. ACG reduces JR and improves View-CS across all baselines.
@@ 4.2 Ablation Study
To verify the effectiveness of each component in our method, we take the generation of ”a DSLR photo of a corgi puppy” as an example and compare the results with and without our design in Figures 7, 8, and 9.
@@
Corse-to-Fine Optimization. In Figure 8, We compare 3D-rendered images generated from the original prompt, “a DSLR photo of a corgi puppy,” with those from a simplified prompt, “a corgi puppy,” using the same seed. DreamFusion may produce unrealistic 3D shapes when using the original prompt, whereas the simplified prompt can somewhat reduce optimization difficulty, leading to a better initialization in the early stages. This initial simplification establishes a foundation for introducing more complex descriptive terms later in the process. However, despite this improvement, the optimized result from the simplified prompt still displays unintended facial features on the back.
@@
Figure 8: Ablation study on coarse to fine optimization based on prompt. For Dreamfusion [38] with seed=2, the top row uses the original prompt ”a DSLR photo of a corgi puppy,” and the bottom row only uses ”a corgi puppy.”
Figure 9: Ablation study on CLIP Pruning. These pseudo-Ground-Truth images identified by CLIP as front-facing are pruned in the back viewpoint range of [120, 240] degrees. This ensures only back view images are retained, enhancing viewpoint consistency in the generated 3D representations.
@@
Viewpoint Rebalancing. Additianally, we analyzed the impact of CLIP Pruning on the balance of pseudo-GT viewpoint distribution during training. As shown in Figure 10, we expected an even distribution across the three viewpoints (front, side, and back), as we randomly sample the three viewpoints with equal probability. However, without the pruning, the pseudo-GT viewpoint reveals a substantial front-view bias with a ratio of approximately 2:1:1. After applying our pruning technique, we observed a significant improvement in balance, achieving a more equitable distribution of 1.2:1:1. This adjustment aligns the pseudo-GT viewpoint more closely with our expectation and is critical for enhancing model’s capacity to overcome Janus problem.
@@ 5 Conclusion
Future Work. Having identified the Janus Problem as a generation bias caused by long-tailed distributions, future work could employ strategies that address long-tailed distribution issues to further alleviate the problem.

### v4

# Improving Viewpoint Consistency in 3D Generation via Structure Feature and CLIP Guidance

@@ 1 Introduction
Figure 2: The long-tailed distribution of viewpoints in the diffusion model’s training data biases the generated viewpoints. The left figure shows a random sample of 6,762 images from Laion2B [34], categorized into three object viewpoints using CLIP ViT-H/14 [ilharco_gabriel_2021_5143773]. The diffusion model maps this long-tailed distribution to a Gaussian distribution, which then produces a similarly biased generation distribution via reverse SDE—fundamentally causing incorrect guidance in 3D representations in the Janus Problem.
Previous studies have attributed the Janus Problem to a generative bias inherent in diffusion models [1, 15, 13, 24]. Given an input prompt, diffusion models tend to generate front-facing images of objects, contributing to geometric inconsistencies in the resulting 3D models. Some methods attempt to mitigate this issue by fine-tuning diffusion models on specialized 3D datasets to enhance multi-view generation capabilities [24, 37, 7]. However, directly fine-tuning on a 3D dataset can lead to catastrophic forgetting and overfitting [14, 33, 47], in addition to being time-consuming and computationally intensive. Alternatively, other approaches address the issue by adjusting sampling strategies to achieve more balanced samples [15]. Nevertheless, without precise knowledge of the long-tailed distribution of the data, such strategies rely on diffusion models to estimate the distribution during sampling, which is often inaccurate [48].
@@ 2.1 Text-to-3D Generation
The task of generating 3D content from text can be broadly divided into two categories. The first category includes end-to-end models such as Point-E [28] and Shap-E [17], which generate 3D outputs directly from text. These models offer fast generation but suffer from poor generalization and subpar quality due to the limited availability of 3D datasets [23]. The second category follows a 2D-to-3D approach inspired by DreamFields [16], where CLIP [30] supervises NeRF [27, 45]. However, the limited supervisory effectiveness of CLIP has prevented this approach from gaining widespread adoption. DreamFusion [29] introduced Score Distillation Sampling (SDS), leveraging pre-trained diffusion models to guide NeRF optimization, which inspired numerous follow-up studies.
Subsequent improvements to DreamFusion have focused on three areas: generation quality, speed, and the Janus problem [44, 22, 21, 4, 15, 46, 8, 5, 42]. Efforts to improve quality mainly revolve around optimizing SDS, as seen in works like ProlificDreamer [44] and LucidDreamer [21], which analyze how SDS leads to mode collapse and excessive smoothing. Speed improvements typically involve optimizing the 3D representation and the generation framework. For example, Magic3D [22] introduces a two-stage optimization strategy and incorporates DMTet [36] for fast surface optimization, and DreamGaussian [42] employs explicit representations to accelerate generation by exploiting the fast rendering speed of Gaussian Splatting [18, 43, 20]. The Janus problem, which concerns generating consistent front and back views, is the primary focus of our research. Previous solutions rely on fine-tuning diffusion models to obtain multi-view capabilities [24, 37]. In contrast, our work takes a different approach by analyzing the inherent generative bias of diffusion models. We achieve superior results without compromising the model’s generalization ability.
@@ 3.1 Rethink Janus Problem
Although numerous studies have identified generation bias in diffusion models as a cause of the Janus Problem [24, 1, 13, 15], a comprehensive proof is still lacking, leaving the issue unresolved. This chapter aims to demonstrate that the long-tailed distribution of viewpoints in the training data of diffusion models is the fundamental factor inducing generation bias, and that the resulting imbalance in viewpoint category guidance for 3D models constitutes the essence of the Janus Problem.
Figure 4: Overview of our work. We adjust cross-attention maps in the diffusion model to enhance focus on viewpoint keywords, guiding the model to generate diverse viewpoints. CLIP guidance prunes incorrect viewpoints during denoising based on a loss threshold, ensuring alignment with specified viewpoints. Decompose the original prompt into object and descriptive words, and optimize the object prompt in the early stages to ensure geometric integrity.
@@
which describes the evolution of the Gaussian distribution over time $s$ from 0 to $T$, corresponding to $t$ decreasing from $T$ to 0. This is essentially the reverse of Eq. 5.
@@
In this context, we interpret SDS as a mechanism that uses images generated from diffusion models [21] to guide images rendered by 3D model. The long-tailed distribution of viewpoints in the training data where the majority of images are front-facing introduces bias into the generated images [35], as shown in Figure 2. When the generated images exhibit a similarly long-tailed distribution [12, 9], the guidance becomes imbalanced, causing the 3D object to overfit to images from the highest-density region.
@@ 3.2 Cross-attention Control
During each noise prediction step in the U-Net of the diffusion process, we identify viewpoint-related keywords in the text prompt: terms that specify spatial viewpoint, such as “front”, “side”, or “back”. These keywords are associated with spatial information that influences the viewpoint of the generated image. We dynamically adjust the attention weights of these viewpoint keywords based on the desired rendering perspective. This adjustment is implemented using the following formula:
@@
Figure 6: Qualitative comparison with baseline methods in text-to-3D generation. From top to bottom: Stable-DreamFusion, Magic3D, LucidDreamer, and our ACG-enhanced versions. For each prompt, we show the 0°, 90°, 180,° and 270° views. ACG consistently recovers the missing back geometry while preserving fine texture, and does so without extra optimisation time or VRAM overhead.
@@ 3.3 CLIP Pruning
where $e_{\text{image}}$ and $e_{\text{text}}$ represent the embeddings of $\bar{x}_{0}$ and $P_{v}$, respectively. The dot product measures semantic alignment, and $\sigma$ is normalized to a range between 1-1 and $1$. If $\sigma$ falls below a specified threshold $\tau$, it indicates that the current generation provides incorrect guidance for the 3D model under the specified viewpoint. In such cases, the update step is skipped to prevent the inclusion of misleading guidance that could cause the Janus problem.
@@
CLIP Pruning mitigates generation bias by eliminating guidance with insufficient semantic view alignment, thereby balancing the distribution of pseudo-GT viewpoints during training. This pruning strategy rebalances the generation bias to some extent, leading to a uniform sampling distribution of viewpoints in 3D generation.
@@ 3.4 Coarse-to-Fine Optimization
We first divide the input prompt into an object prompt (main subject) and a description prompt (additional characteristics). In the initial stage, we focus on the object prompt by adjusting cross-attention maps and applying CLIP pruning to amplify attention to spatial viewpoint terms. This ensures the generated 3D model accurately captures the core geometry and specified viewpoint.
@@ 4.1 Text-to-3D Generation
Quantitative Comparison. Previous research in 3D generation lacks standardized metrics. Existing works typically use 2D image evaluation metrics to assess the quality or prompt relevance of generated content. However, these metrics are insufficient for evaluating the Janus problem. To assess the consistency of 3D geometry, we generate 3D objects using 20 randomly selected text prompts from the DreamFusion prompt dataset. For each 3D object, we count the instances of inconsistent 3D content produced by each approach, defining this count as the Janus Problem Rate (JR). As shown in Table 1, our method significantly reduces the multi-head problem.
@@
LucidDreamer [21] 80.00
LucidDreamer+ACG (Ours) 30.00
Magic3D [22] 75.00
Magic3D+ACG (Ours) 35.00
DreamFusion [29] 80.00
DreamFusion+ACG (Ours) 35.00
Table 1: Comparison of geometric consistency between baseline methods and ACG-enhanced versions. Lower JR indicates fewer Janus outcomes. ACG reduces JR across all baselines.
@@ 4.2 Ablation Study
To verify the effectiveness of each component in our method, we take the generation of “a DSLR photo of a corgi puppy” as an example and compare the results with and without our design in Figures 7, 8, and 9.
@@
Coarse-to-Fine Optimization. In Figure 8, we compare 3D-rendered images generated from the original prompt, “a DSLR photo of a corgi puppy”, with those from a simplified prompt, “a corgi puppy”, using the same seed. DreamFusion may produce unrealistic 3D shapes when using the original prompt, whereas the simplified prompt can somewhat reduce optimization difficulty, leading to a better initialization in the early stages. This initial simplification establishes a foundation for introducing more complex descriptive terms later in the process. However, despite this improvement, the optimized result from the simplified prompt still displays unintended facial features on the back.
@@
Figure 8: Ablation study on coarse to fine optimization based on prompt. For Dreamfusion [41] with seed=2, the top row uses the original prompt “a DSLR photo of a corgi puppy”, and the bottom row only uses “a corgi puppy”.
Figure 9: Ablation study on CLIP Pruning. These pseudo-Ground-Truth images identified by CLIP as front-facing are pruned in the back viewpoint range of [120, 240] degrees. This ensures only back-view images are retained, enhancing viewpoint consistency in the generated 3D representations.
@@
Viewpoint Rebalancing. Additionally, we analyzed the impact of CLIP Pruning on the balance of pseudo-GT viewpoint distribution during training. As shown in Figure 10, we expected an even distribution across the three viewpoints (front, side, and back), as we randomly sample the three viewpoints with equal probability. However, without the pruning, the pseudo-GT viewpoint reveals a substantial front-view bias with a ratio of approximately 2:1:1. After applying our pruning technique, we observed a significant improvement in balance, achieving a more equitable distribution of 1.2:1:1. This adjustment aligns the pseudo-GT viewpoint more closely with our expectation and is critical for enhancing model’s capacity to overcome Janus problem.
@@ 5 Conclusion
Future Work. Having identified the Janus Problem as a generation bias caused by long-tailed distributions, future work could pursue two complementary directions: (i) fine-tuning the CLIP on the noise-corrupted renderings that arise during diffusion, so its similarity signals remain balanced across both frequent and rare viewpoints, and (ii) reshaping the SDS guidance by injecting tail-view exemplars, explicitly drawing the optimisation gradients toward the low-density regions of viewpoint space.